OpenStudy (anonymous):

The formula below gives V, the volume of a cylinder with radius r and height 16 V(r) = 16r2 What is the value of V(5)?

OpenStudy (anonymous):

Plug in the value of 5 for every r value of the functions, so: V(5) = 16*5*2 = 160 :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Replace 5 for r, V(5) = 16*5*2 V(5) = 160

OpenStudy (anonymous):

Hmph what grade is this for ^~^

OpenStudy (anonymous):

9th cx

OpenStudy (anonymous):

;O wuuuut lol Gl

OpenStudy (anonymous):

why is it surprising ? XDD